OpenStudy (anonymous):

Find the annual interest rate if a $5000 investment earns $550 annual simple interest in one year.

OpenStudy (anonymous):

I = prt 1) nope: since she earned 1536, that is your interest (I) 1536 = p*0.04*8 1536 = 0.32p 1536/0.32 = 0.32p/0.32 p = $4800 2) 550 = 5000*r*1 550/5000 = 5000r/5000 0.11 = r = 11%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

UR welcome